# Multi-messenger signatures of a deformed magnetar in gamma-ray bursts

### v1

@@ 3.1 Dipole luminosity
Energy loss via EM dipole moment (dipole luminosity) is given by:
@@
where $P=2\pi/\Omega$, $B_{\rm dip}$ is the dipole magnetic field and $\alpha$ is the angle between the magnetic dipole moment and the rotational axis. hereafter, we select $\alpha=\pi/2$. In Eq. (16) we use appropriate values for scaling parameters.Combining Eqs. (16) and (7) the following equation for the angular momentum evolution is evaluated:
@@
Finally putting Eqs. (5), (9) and (17) toghether, evolutionary equation for eccentricity in incompressible fluid is:
@@
On the other hand, by using (5), (15) and (17) for the compressible fluid can obtain :
@@
3.2 Quadrapole luminosity
Energy loss for EM quadrapole magnetic moment (quadrupole luminosity) is given by:
@@ 3.2 Quadrapole luminosity
where $B_{\rm quad}$ is the quadrapole magnetic field and $\beta_{1}$ and $\beta_{2}$ are the inclination angles of the magnetic moments. It is instructive to introduce $\kappa$ as the quadrupole to dipole ratio:
@@
An order of magnitude estimation for EM quadrapole luminosity is given by:
@@
$=5.08\times 10^{51}erg/s\left(\frac{B}{10^{15}G}\right)^{2}\left(\frac{a_{0}}{10^{6}cm}\right)^{8}\left(\frac{0.5\rm ms}{P}\right)^{4}\left(\frac{\kappa}{500}\right).$
Therefore, angular momentum loss due to quadrapole magnetic moment radiations is:
@@
where $\xi=-32/135B_{\rm quad}^{2}a_{0}^{8}\kappa$, leading to the evolutionary equation for the eccentricity of the incompressible model as:
@@
For compressible model this evolutionary equation is evaluated as:
@@ 3.3 Total EM Luminosity
Now, we consider the total EM radiation in order to determine the angular momentum loss which is expressed as:
@@
Therefore, the following equations describe the evolution of eccentricity for the incompressible fluid as :
@@
and in the case of the compressible model we obtain :
@@
Figure 2: Total EM luminosity for the incompressible model with fixed mass $M_{T}=2M\odot$. Different magnetic fields and quadrupole to dipole ratios are labeled.
@@
Figure 3: EM dipole luminosity for the incompressible and compressible models, Here we assume that $M_{T}=2M\odot$ and $B=10^{15}G$. In the case of the compressible model, different adiabatic indexes are labeled.
@@
Figure 4: EM dipole luminosity for the compressible model is plotted for $\Gamma=1.43$, where magnetic field and mass of the magnetar are labeled.
@@
Figure 5: EM dipole and quadrupole emission of the compressible model with $\Gamma=1.43$, $B=10^{16}G$ and different fraction of dipole to quadrupole moments as labeled.
@@
Fig. 2 shows the total EM luminosity for a magnetar containing incompressible fluid. Here, we assume that the dipole magnetic field strength is $B=10^{14}-10^{16}\rm G$ and the quadrupole is up to $500$ times larger than the dipole moment. It is seen that for higher ratios of magnetic quadrupole to dipole moments $\kappa$, the EM light curve of the magnetar starts to decay at earlier times with lower slope. This is a result of spin down with a faster rate at earlier times. We find that the luminosity is mainly powered by the dipole magnetic field at the earlier times and at later times it is dominated by dipole emission.
@@
In Fig. 3, the effect of different polytropic equations of states (EoSs) on the dipole EM luminosity is investigated. It is illustrated that for some specific values of adiabatic index $\Gamma$, close enough to 4/3, a peak in the light curve appears. As it is seen, the peak fades away as $\Gamma$ increases and goes beyond $4/3$. These regions within the parameter space where the spin-up is caused by the decreasing moment of inertia due to the softness of the EoS, and black hole formation does not occur, we attribute the resulting flaring behavior in the light curve to these regions. The flaring behavior is closely tied to the rotational dynamics of new born NS which leads to a temporary spin-up of NS during its early spin-down phase. This effect crucially depends on the applied EoS and the adiabatic index $\Gamma$, as it is seen from Eq. (19) the non-axisymmetric deformations of the NS including a compressible fluid for $\Gamma\approx 4/3$ will show a non trivial behavior leading to a temporary spin-up behavior that are reflected in the observed flares upon the light curves.
@@
The incompressible model is parameterized with $\Gamma=\infty$ and as it is seen from Fig. 3, it has the highest maximum initial luminosity for a given magnetic field and mass of the magnetar, and the maximum initial luminosity increases as softer EoSs are applied. Additionally, one can see that for stiffer EOSs, the decay time in the light curve occurs in the earlier times. Note that, the late time behavior of the light curve for all models with different $\Gamma$ values is similar. In Fig. 4, we explore the effect of both mass and magnetic field on the light curve of the dipole EM radiation for a compressible fluid with $\Gamma=1.43$. As shown in Fig. 3, the magnetic field has a significant impact on the peak luminosity, $L_{\rm p}$ and the peak time, $T_{\rm p}$, while the mass’s effect in a given range on $L_{\rm p}$ and $T_{\rm p}$ is not significant. The allowed range of the parameter space of mass and $\Gamma$ is determined through the stability analysis of the magnetar under deformation (see Appendix (A)). We introduce the valid mass range for our model based on black hole formation criteria following the method presented in Finn & Shapiro (1990); Wheeler & Klauder (1972)
@@
In Fig. 5, the magnetic dipole luminosity is compared to the total EM luminosity including quadrupole emission with different $\kappa$ in the compressible model. We find that taking the quadrupole magnetic moment into account lead to an earlier peak time in the light curves. Moreover, for a given dipole magnetic field, the effect of the quadrupole to dipole moment ratio on the peak luminosity is negligible.
4 GW radiation and Spin down
@@ 4.1 GW luminosity
In this section, we investigate the GW radiations from a deformed magnetar due to the asymmetry in its equatorial plane. The GW luminosity of a deformed magnetar is expressed as Andersson (2019):
@@
$=1.80\times 10^{53}\,\text{erg/s}\left(\frac{a_{0}}{10^{6}\,\text{cm}}\right)^{4}\left(\frac{\epsilon}{10^{-3}}\right)^{2}\left(\frac{0.5\,\text{ms}}{P}\right)^{6}\left(\frac{M}{2.6M_{\odot}}\right)^{2}.$ (30)
By introducing $\gamma_{0}\equiv 32I_{0}^{2}\epsilon^{2}/5$, the time evolution of the eccentricity due to GW radiation from an incompressible fluid is given by the following equation :
@@
Figure 6: GW luminosity for compressible model for $\Gamma=1.43$, and $\epsilon=10^{-3}$ for different masses of NS as labeled.
Figure 7: GW luminosity for compressible model for $\Gamma=1.53$, and $\epsilon=10^{-3}$ for different masses of NS as labeled.
Figure 8: GW luminosity for incompressible ($\Gamma=\infty$) and compressible models, Here we assume that $M_{T}=2M\odot$ and $\epsilon=10^{-3}$, in the case of compressible model different adiabatic indexes ($\Gamma$) are labeled.
We plot $L_{GW}$ for some cases in Fig. 6 - 8 to examine the gravitational energy loss from a newborn magnetar. Fig. 6 and Fig. 7 describe GW luminosity for the compressible model for $\Gamma=1.43$ and $\Gamma=1.53$ respectively.These light curves indicate that for higher masses, the luminosity peak appears at earlier times with higher magnitude. At late times light curves for all masses coincide. Comparing Fig. 6 and Fig. 7, light curve peaks in Fig. 6 are more apparent. As a result, a higher frequency signal should be detected for a magnetar with softer EOS during peak time. Fig. ‘8 demonstrates GW luminosity for different EOSs. As $\Gamma$ tends to $4/3\approx 1.33$, the peak in the light curves fades away, which is the same result as seen in the EM light curve. Furthermore, the behavior of maximum luminosity with respect to changes in $\Gamma$ is the same.
@@ 4.2 Detectability of the GW Signals
Here, we consider the detectiblity of the GW signal originating from a deformed magnetar in the light of future GW detectors. The GW strain for the deformed magnetar located at distance d from the observer is given by Howell et al. (2011); Bambi et al. (2021); Corsi & Mészáros (2009):
@@
Knowing that $f=\Omega/\pi$ one can find that:
@@
In order to find the explicit form of $h_{c}$ one needs to find the term $dt/df$, which can be find according to:
@@
Considering the domination of GW radiation at early times:
@@
which leads to the following characteristic GW strain as:
@@
In order to find $h(t)$ one needs to determine $\Omega(t)$ due to variation of eccentricity, we plot $h_{c}$ in Fig. 9 for different masses as labled assuming $d=40Mpc$, $a_{0}=10^{6}cm$ and $P=0.5$ ms for a new born magnetar. It is seen than the variation of the mass in the given range has not significant impact on $h_{c}$ and the resulting curves for three different values of masses almost coincide each other.
On the other hand, for later times, EM radiation dominates the energy loss, by rewriting Eqs. (36) and (37) for EM case, $dt/df$ is given as bellow:
$\frac{dt}{df}\approx\frac{\pi^{2}If}{L_{EM}},$ (39)
the characteristic GW strain for this regime is obtained as:
@@
Here, we can ignore the effect of quadruple magnetic field while $\kappa a^{2}\Omega^{2}<<1$, and write $h_{c}(t)$ only for EM dipole radiations:
@@
Characteristic GW strain for the case in which the spin down drives by the EM dipole radiation is plotted in Fig. 9 for various $B$ and $\epsilon$ as labeled. We see from Eq. (41) and Fig. 9 that a large ellipticity would produce a stronger gravitational wave signal. Eq. (41) is inversely proportional to the $B_{\text{dip}}$, however the magnetic field should be large enough to reproduce the observed EM luminosity therefore magnetic field strength has a lower bound. On the other hand it was shown that the large ellipticity can be generated thanks to high values of the magnetic field, which means larger magnetic field components Xie et al. (2022b); Sur & Haskell (2021); Lasky & Glampedakis (2016); Stella et al. (2005).
The projected sensitivities for aLIGO O5, Einstein Telescope (ET) and Cosmic Explorer (CE) detectors are also presented in Fig. 9, it is found that GW signals from magnetars with the specific set of parameters, reach the sensitivity threshold of theses detectors. However, being $h_{c}$ within the sensitivity range of GW detector (as it is seen from Fig. 9) does not necessarily guarantee that the GW will be detected, though it suggests that detection should be possible under ideal conditions. In practice, the detector’s ability to identify a GW signal depends on several factors, such as noise, duration of the signal, and the methods used to search for the signal. If the signal to noise ratio (SNR) is too low (e.g., the signal is weak or buried in background noise), it may not be detectable. For continuous or long-duration signals (such as those expected from deformed neutron stars or magnetars), noise can accumulate over time, and the signal can be buried in background noise, making detection harder despite being within the sensitivity range. Imperfect search algorithms or unknown signal characteristics could also result in a missed detection. Therefore, being in the sensitivity range is a necessary but not sufficient condition for the detection. However, if the GW signal associated with the EM light curve of the magnetar can be detected in the future, it would be a strong evidence that the magnetar can act as the central engine of GRBs.
@@
Figure 9: Sensitivities of GW detectors to GW characteristic strain produced by rotating magnetars for different masses as labeled.
[!htbp]
5 GW and EM spin down

### v2

@@ 3.1 Dipole luminosity
Energy loss via EM dipole moment (dipole luminosity) is given by
@@
where $P=2\pi/\Omega$, $B_{\rm dip}$ is the dipole magnetic field and $\alpha$ is the angle between the magnetic dipole moment and the rotational axis. Hereafter, we select $\alpha=\pi/2$. In Eq. (18) we use appropriate values for scaling parameters. Combining Eqs (18) and (7) the following equation for the angular momentum evolution is evaluated:
@@
Finally combining Eqs (5), (10) and (19), evolutionary equation for eccentricity in incompressible fluid is
@@
On the other hand, by using Eqs (5), (17) and (19) for the compressible fluid can be obtain as
@@
3.2 Quadrupole Luminosity
Energy loss for EM quadrupole magnetic moment (quadrupole luminosity) is given by
@@ 3.2 Quadrupole Luminosity
where $B_{\rm quad}$ is the quadrupole magnetic field and $\beta_{1}$ and $\beta_{2}$ are the inclination angles of the magnetic moments. It is instructive to introduce $\kappa$ as the quadrupole-to-dipole ratio:
@@
An order of magnitude estimation for EM quadrupole luminosity is given by
@@
$=5.08\times 10^{51}erg/s\left(\frac{B}{10^{15}G}\right)^{2}\left(\frac{a_{0}}{10^{6}cm}\right)^{8}\left(\frac{0.5\rm ms}{P}\right)^{6}\left(\frac{\kappa}{500}\right).$
Therefore, angular momentum loss due to quadrupole magnetic moment radiations is
@@
where $\xi=-32/135B_{\rm quad}^{2}a_{0}^{8}\kappa$, leading to the evolutionary equation for the eccentricity of the incompressible model as
@@
For the compressible model, this evolutionary equation is evaluated as
@@ 3.3 Total EM Luminosity
Now, we consider the total EM radiation in order to determine the angular momentum loss, which is expressed as
@@
Therefore, the following equations describe the evolution of eccentricity for the incompressible fluid as
@@
and in the case of the compressible model, we obtain
@@
Figure 2: Total EM luminosity for the incompressible model with fixed mass $M_{T}=2M\odot$. Different magnetic fields and quadrupole-to-dipole ratios are labeled.
@@
Figure 3: EM dipole luminosity for the incompressible and compressible models. Here, we assume that $M_{T}=2M\odot$ and $B=10^{15}G$. In the case of the compressible model, different adiabatic indexes are labeled.
@@
Figure 4: EM dipole luminosity for the compressible model is plotted for $\Gamma=1.43$, where the magnetic field and mass of the magnetars are labeled.
@@
Figure 5: EM dipole and quadrupole emission of the compressible model with $\Gamma=1.43$, $B=10^{16}G$ and different fractions of dipole to quadrupole moments as labeled.
@@
Fig. 2 shows the total EM luminosity for a magnetar containing incompressible fluid. Here, we assume that the dipole magnetic field strength is $B=10^{14}-10^{16}\rm G$ and the quadrupole is up to $500$ times larger than the dipole moment. It is seen that for higher ratios of magnetic quadrupole-to-dipole moments $\kappa$, the EM light curve of the magnetar starts to decay at earlier times with a lower slope. This is a result of spin-down with a faster rate at earlier times. We find that the luminosity is mainly powered by the dipole magnetic field at the earlier times, and at later times, it is dominated by dipole emission.
@@
In Fig. 3, the effect of different polytropic EoSs on the dipole EM luminosity is investigated. It is illustrated that for some specific values of adiabatic index $\Gamma$, close enough to 4/3, a peak in the light curve appears. As can be seen, the peak fades away as $\Gamma$ increases and goes beyond $4/3$. These regions within the parameter space where the spin-up is caused by the decreasing moment of inertia due to the softness of the EoS, and BH formation does not occur, we attribute the resulting flaring behavior in the light curve to these regions. The flaring behavior is closely tied to the rotational dynamics of new born NS, which leads to a temporary spin-up of the NS during its early spin-down phase. This effect crucially depends on the applied EoS and the adiabatic index $\Gamma$, as can be seen from Eq. (21). The non-axisymmetric deformations of the NS, including a compressible fluid for $\Gamma\approx 4/3$ will show a nontrivial behavior leading to a temporary spin-up behavior that is reflected in the observed flares upon the light curves.
@@
The incompressible model is parameterized with $\Gamma=\infty$ and as it is seen in Fig. 3, it has the highest maximum initial luminosity for a given magnetic field and mass of the magnetar, and the maximum initial luminosity increases as softer EoSs are applied. Additionally, one can see that for stiffer EoSs, the decay time in the light curve occurs at earlier times. Note that the late-time behavior of the light curve for all models with different $\Gamma$ values is similar. In Fig. 4, we explore the effect of both mass and magnetic field on the light curve of the dipole EM radiation for a compressible fluid with $\Gamma=1.43$. As shown in Fig. 3, the magnetic field has a significant impact on the peak luminosity, $L_{\rm p}$ and the peak time, $T_{\rm p}$, while the mass’s effect in a given range on $L_{\rm p}$ and $T_{\rm p}$ is not significant. The allowed range of the parameter space of mass and $\Gamma$ is determined through the stability analysis of the magnetar under deformation (see Appendix (A)). We introduce the valid mass range for our model based on black hole formation criteria following the method presented in Finn & Shapiro (1990); Wheeler & Klauder (1972)
@@
In Fig. 5, the magnetic dipole luminosity is compared to the total EM luminosity, including quadrupole emission with different $\kappa$ in the compressible model. We find that taking the quadrupole magnetic moment into account leads to an earlier peak time in the light curves. Moreover, for a given dipole magnetic field, the effect of the quadrupole-to-dipole moment ratio on the peak luminosity is negligible.
4 GW Radiation and Spin-down
@@ 4.1 GW Luminosity
In this section, we investigate the GW radiation from a deformed magnetar due to the asymmetry in its equatorial plane. The GW luminosity of a deformed magnetar is expressed as Andersson (2019)
@@
$=1.73\times 10^{50}\,\text{erg/s}\left(\frac{a_{0}}{10^{6}\,\text{cm}}\right)^{4}\left(\frac{\epsilon}{10^{-3}}\right)^{2}\left(\frac{0.5\,\text{ms}}{P}\right)^{6}\left(\frac{M}{2M_{\odot}}\right)^{2}.$ (32)
By introducing $\gamma_{0}\equiv 32I_{0}^{2}\epsilon^{2}/5$, the time evolution of the eccentricity due to GW radiation from an incompressible fluid is given by the following equation
@@
Figure 6: GW luminosity for the compressible model with $\Gamma=1.43$, and $\epsilon=10^{-3}$ for different masses of NSs as labeled.
Figure 7: GW luminosity for the compressible model with $\Gamma=1.53$, and $\epsilon=10^{-3}$ for different masses of NSs as labeled.
Figure 8: GW luminosity for the incompressible ($\Gamma=\infty$) and compressible models. Here, we assume that $M_{T}=2M\odot$ and $\epsilon=10^{-3}$, in the case of the compressible model, different adiabatic indexes ($\Gamma$) are labeled.
We plot $L_{GW}$ for some cases in Fig. 6 - 8 to examine the gravitational energy loss from a newborn magnetar. Fig. 6 and Fig. 7 describe GW luminosity for the compressible model for $\Gamma=1.43$ and $\Gamma=1.53$, respectively. These figures indicate that for higher masses, the luminosity peak appears at earlier times with a higher magnitude. At late times, luminosity curves for all masses coincide. It is found that a higher amplitude signal might be detected for a magnetar with softer EoS during peak time. Fig. 8 illustrates the GW luminosity for different EoSs. As $\Gamma$ tends to large values far beyond $4/3\approx 1.33$, the peak in $L_{GW}$ fades away, which is the same result as seen in the EM light curve. Furthermore, the behavior of maximum luminosity with respect to changes in $\Gamma$ is similar.
@@ 4.2 Detectability of the GW Signals
Here, we consider the detectability of the GW signal originating from a deformed magnetar in the light of future GW detectors. The GW strain for the deformed magnetar located at distance D from the observer is given by Howell et al. (2011); Bambi et al. (2021); Corsi & Mészáros (2009)
@@
Knowing that $f=\Omega/\pi$ one can find that
@@
In order to find the explicit form of $h_{c}$ one needs to find the term $dt/df$, which can be find according to
@@
Considering the domination of GW radiation at early times,
@@
which leads to the following characteristic GW strain:
@@
In order to find $h(t)$ one needs to determine $\Omega(t)$ due to variation of eccentricity, we plot $h_{c}$ in Fig. 9 for different masses as labeled assuming $D=40\text{Mpc}$, $a_{0}=10^{6}cm$ and $P=0.5$ ms for a new born magnetar. It is seen that the variation in the mass in the given range has no significant impact on $h_{c}$ and the resulting curves for the three different values of masses almost coincide with each other.
On the other hand, for later times, EM radiation dominates the energy loss, by rewriting Eqs. (38) and (45) for the EM case, $dt/df$ is given as bellow:
$\frac{dt}{df}\approx\frac{\pi^{2}If}{L_{EM}}.$ (41)
The characteristic GW strain for this regime is obtained as
@@
Here, we can ignore the effect of quadruple magnetic field, while $\kappa a^{2}\Omega^{2}<<1$, and write $h_{c}(t)$ only for the EM dipole radiation:
@@
The characteristic GW strain for the case in which the spin-down is driven by the EM dipole radiation is plotted in Fig. 9 for various $B$ and $\epsilon$ as labeled. We see from Eq. (43) and Fig. 9 that a large ellipticity would produce a stronger GW signal. Eq. (43) is inversely proportional to the $B_{\text{dip}}$, however, the magnetic field should be large enough to reproduce the observed EM luminosity, therefore magnetic field strength has a lower bound. On the other hand, it has been shown that the large ellipticity can be generated thanks to high values of the magnetic field, which means larger magnetic field components Xie et al. (2022b); Sur & Haskell (2021); Lasky & Glampedakis (2016); Stella et al. (2005).
The projected sensitivities for the ALIGO O4, ALIGO O5, KAGRA O5, Einstein Telescope (ET), and Cosmic Explorer (CE) detectors are also presented in Fig. 9. It is found that GW signals from magnetars with a specific set of parameters reach the sensitivity threshold of these detectors. Here, we report a new signature of a new born magnetar in the GW strain, where there is a turning point at high frequencies due to the spin-up and spin-down for a range of compressible EoSs (here, we select $\Gamma=1.43$). This feature may be used to discriminate the GW signal generated by NS from those produced by BH central engine van Putten et al. (2024). We will consider this case in more detail in a follow-up paper.
In order to consider the detectability of the signal, we also calculate the root sum squared strain for GW as below (Abbott et al. (2017)):
[…]
where $\tilde{h}_{+}(f)$ and $\tilde{h}_{\times}(f)$ are the plus and cross polarizations of GW, respectively. For the case of a deformed magnetar, we use the relations described in Finn & Shapiro (1990) for $\tilde{h}_{+}$ and $\tilde{h}_{\times}$ polarizations, assuming the same contribution of both polarizations. The frequency range is evaluated from Eq. (3), $f=2\Omega$, for the compressible model to be $f_{min}=103Hz$ and $f_{max}=6994Hz$, which corresponds to the eccentricity changing from 0.81 to 0.005. Therefore, we have
$h_{rss}=$ $1.88\times 10^{-23}Hz^{-1/2}$ (45)
$\left(\frac{\epsilon}{10^{-3}}\right)\left(\frac{40\text{Mpc}}{\text{D}}\right)\left(\frac{\text{M}}{2M_{{}_{\odot}}}\right)^{1/2}\left(\frac{R}{10\ \text{km}}\right).$
The signal-to-noise ratio (SNR) for a narrowband source at the central frequency $f_{0}$ is given by
$\frac{S}{N}=\Theta\frac{h_{rss}}{\sqrt{S(f_{0})}}$ (46)
where $\Theta$ is the angle factor and for an optimally oriented source can be estimated as $\Theta_{rms}=\sqrt{2/5}$ Sutton (2013). Considering $S(f_{0}\approx 1\text{ kHz})$, $h_{rss}$ is calculated for a given set of scaling parameters as in Eq. (45), and the SNRs for the current and future detectors are presented in Table 1. This result shows that GW detectors might be able to detect GWs from newborn magnetars and record the specific signatures of spin evolution and deformation of the star.
Table 1: Sensitivity and SNR values for different detectors.
GW Detector $S(f_{0}\approx 1\text{ kHz})$$/\sqrt{Hz}$ SNR
ALIGO O4 $5.75\times 10^{-24}$ 2.02
KAGRA O5 $1.28\times 10^{-23}$ 0.93
ALIGO O5 $2.53\times 10^{-24}$ 4.69
ET $5.76\times 10^{-25}$ 20.62
CE $3.27\times 10^{-25}$ 36.32
@@
Figure 9: Sensitivities of GW detectors to GW characteristic strain produced by rotating magnetars for different masses, as labeled.
Note that being the GW strain within the sensitivity range of the GW detector (as it is seen from Fig. 9) does not necessarily guarantee that the GW will be detected, though it suggests that detection should be possible under ideal conditions. In practice, the detector’s ability to identify a GW signal depends on several factors, such as noise, duration of the signal, and the methods used to search for the signal. If the SNR is too low (e.g., the signal is weak or buried in background noise), it may not be detectable. For continuous or long-duration signals (such as those expected from deformed NSs or magnetars), noise can accumulate over time, and the signal can be buried in background noise, making detection harder despite being within the sensitivity range. Imperfect search algorithms or unknown signal characteristics could also result in a missed detection. Therefore, being in the sensitivity range is a necessary but not sufficient condition for the detection. However, if the GW signal associated with the EM light curve of the magnetar can be detected in the future, it would be a strong evidence that the magnetar can act as the central engine of GRBs.
5 GW and EM spin-down